OpenStudy (blaque678):

The sum of two numbers is 48.One number is 3 times as large as the other. What are the numbers?

OpenStudy (shane_b):

x+y=48 x=3y Now just solve for x and y (I'd use the substitution method for this one)

OpenStudy (blaque678):

thats where i get confused

OpenStudy (tyteen4a03):

Which part are you confused? By substitution you substitute x=3y into x+y=48, which will become 3y+y = 48. Solve this equation to get the value of y, then use y to get x.

OpenStudy (blaque678):

when it comes to breaking it down and subsitituing

OpenStudy (shane_b):

You have 2 equations with 2 unknown values: x+y=48 x=3y Using the substitution method you'd take the second equation which tells you exactly what x equals and plug it into the first equation anywhere you see an x: (3y)+y=48 4y=48 y=12 Now you know what y is. Use either of the original equations plugging in y to solve for x: x + y = 48 x+12=48 (plug in y=12) x=48-12=36 Therefore, x must equal 36 and y = 12.